One grandfather
One father
One grandson
The 3 together are 140 years old
The grandfather's age is equal to the number of months of the grandson
The father's age in weeks is equivalent to the grandsons age in days
Tell me what are the individual ages of the 3?

Found 2 solutions by sachi, richwmiller:
Answer by sachi(548)

let grandson is x yrs old=12x months=12x*30 days=360x days
The grandfather's age is equal to the number of months of the grandson
so age of grandfather=12x yrs
The father's age in weeks is equivalent to the grandsons age in days
so father's age in weeks=360x
so father's age in yrs=360x/52=90x/13
The 3 together are 140 years old
so x+12x+90x/13=140
or 13x+156x+90x=140*13=1820
or 259x=1820
or x=7.02=7 yrs
so age of grandfather=12x yrs=12*7=84 yrs
so father's age in yrs=360x/52=90x/13=90*7/13=48.46=49 yrs
the individual ages of the 3 are 7,49,84
we have taken 1 yr=12 months
1 month =30 days
1yr=52weeks

Answer by richwmiller(17219)

g+f+s=140,
12s=g,
364s=52f

f = 49, g = 84, s = 7
The hard part was deciding what to use for number of days. I decided 52*7=364
which would make the father 7 times the son